# README: Bridging Medical Jargon and Lay Understanding for Patient Education through Data-Centric NLP

### v2

@@ Abstract
The advancement in healthcare has shifted focus toward patient-centric approaches, particularly in self-care and patient education, facilitated by access to Electronic Health Records (EHR). However, medical jargon in EHRs poses significant challenges in patient comprehension. To address this, we introduce a new task of automatically generating lay definitions, aiming to simplify complex medical terms into patient-friendly lay language. We first created the README dataset, an extensive collection of over 50,000 unique (medical term, lay definition) pairs and 300,000 mentions, each offering context-aware lay definitions manually annotated by domain experts. We have also engineered a data-centric Human-AI pipeline that synergizes data filtering, augmentation, and selection to improve data quality. We then used README as the training data for models and leveraged a Retrieval-Augmented Generation method to reduce hallucinations and improve the quality of model outputs. Our extensive automatic and human evaluations demonstrate that open-source mobile-friendly models, when fine-tuned with high-quality data, are capable of matching or even surpassing the performance of state-of-the-art closed-source large language models like ChatGPT. This research represents a significant stride in closing the knowledge gap in patient education and advancing patient-centric healthcare solutions   [^1] Our code and data will be released at https://github.com/seasonyao/NoteAid-README.
@@ 1 Introduction
The advancement in healthcare transcends medical breakthroughs, encompassing an increased emphasis on patient involvement in self-care Boling (2009); Spooner et al. (2019); Baldry et al. (1986); Schillinger et al. (2009). The criticality of effective patient education is underscored in an era marked by the emergence of new treatments that demand patient comprehension and cooperation. Access to Electronic Health Records (EHR) has become a foundational element of modern educational paradigms Gabay (2017); Lye et al. (2018). Despite advancements in electronic medical record management Walker et al. (2019), one significant barrier persists in the form of medical jargon   [^2] A “medical jargon term” is the specialized language healthcare professionals use that can be complex for non-medical individuals. A “lay definition” translates this jargon into accessible language for the general public, aiming to bridge the understanding gap and enhance the communication of health-related information. in EHRs, impeding patient understanding and self-care Kujala et al. (2022); Choudhry et al. (2016); Khasawneh et al. (2022). As shown in Figure 1, tools like NoteAid Chen et al. (2018); Kwon et al. (2022), which employs Natural Language Processing (NLP) to demystify complex medical terms, have been instrumental in bridging the communication gap between healthcare professionals and patients Lalor et al. (2018, 2021, 2023). However, lay language dictionary resources like the Consumer Health Vocabulary (CHV) Zeng and Tse (2006); He et al. (2017); Ibrahim et al. (2020) are limited in scale, posing a challenge to NoteAid. For instance, only a fraction (about 4%) of the medical terms in NoteAid have been annotated with lay definitions, highlighting the need for a more scalable solution to address this knowledge gap effectively.
@@
Addressing this issue requires shifting our focus to online health education materials such as Unified Medical Language System (UMLS)  Bodenreider (2004), MedlinePlus Patrias and Wendling (2007), Wikipedia, and Google. However, as indicated in Appendix Table 4, these resources often present too difficult information for the average patient to understand. For example, these resources’ average readability measured by the Flesch Kincaid Grade Level Solnyshkina et al. (2017) was post-secondary or higher education, while the average readability of a US adult was 7-8th grade level Doak et al. (1996, 1998); Eltorai et al. (2014). To bridge this gap, we have engaged medical experts to meticulously curate lay definitions for jargon terms found in NoteAid-MedJEx Kwon et al. (2022), targeting a comprehension level suitable for individuals with a 7th to 8th-grade education. Each jargon term has been redefined across various contexts, ensuring their applicability in diverse clinical scenarios. This effort led to our creation of the REsource of lAy Definitions for MEdical jargon (README) dataset, an expansive resource containing over 51,623 (medical jargon term, lay definition) pairs. The README dataset comprises an impressive 308,242 data points, each consisting of a clinical note context, a medical jargon term, and its corresponding lay definition. Thus, the dataset significantly enhances the accessibility and comprehensibility of medical information for patients.
@@
  • Introduced a new task of automatically generating lay definitions for medical jargon. We created a substantial expert-annotated dataset of 308,242 data points that can be used directly as a detailed lay-language dictionary for patient education tools such as NoteAid and as training data for this new task. We leveraged a Retrieval-Augmented Generation (RAG) method to reduce hallucinations and improve the quality of model outputs. Our work aims to enhance patient education by improving patients’ comprehension of medical documentation.
@@ 3.2 Lay Definition Annotation
Domain-experts read each sentence and identified as medical jargon terms that would be considered difficult to comprehend for anyone no greater than a 7th-grade education  [^3] The rule of thumb is that if a candidate term has a lay definition comprehensible to a 4-7th grader as judged by Flesch-Kincaid Grade Level (Solnyshkina et al., 2017), the candidate term is included as a jargon term.. Overall, 51,623 unique (medical jargon term, lay definition) pairs with 308,242 mentions  [^4] So there are 308,242 data points with (EHR context, medical jargon term, lay definition) format. in the EHR repository have been annotated by complying with the annotation guidelines presented in Appendix A.
@@ 3.3 General Definition Retrieval
We then employed the Scispacy library  [^5] We used the Scispacy en_core_sci_lg model to obtain the data. https://github.com/allenai/scispacy to retrieve the corresponding UMLS definitions Lindberg et al. (1993) of these annotated medical jargon terms as the general definitions. We follow the retrieval and preprocessing steps in Appendix C to filter the valid general definitions for the README dataset. This preliminary cleaning results in 308,242 data points in the README-exp (e.g., expert annotated) dataset, each consisting of a clinical note context, a medical jargon term, a corresponding lay definition, and a corresponding general definition, as shown in Figure 2.
@@ 3.4 Examiner-Augmenter-Examiner (EAE)
3.4.1 Examiner (expert-annotated data)
Initially, basic data cleaning, as outlined in Appendix C, was applied. To enhance this, we employed GPT-3.5-turbo, using a few-shot learning approach with seven examples - four demonstrating acceptable data points and three showing unacceptable ones. These prompts served as the ‘Human’ element in our Human-AI-in-the-loop model, as depicted in Figure 2 and detailed in Algorithm 1. The prompts are detailed in Table 6. We choose GPT-3.5-turbo here because our evaluation (Section 4.3 and Appendix D) shows that the definitions it generates for medical terms can reach a human-satisfying level. Post-cleaning  [^6] More GPT Running Details in Examiner step can be found on Appendix B.4, approximately 39% of UMLS general definitions were deemed suitable by the Examiner (e.g., GPT-3.5-turbo). The suitable 113,659 data points were archived in R-exp_good, while the unsuitable 177,140 ones were stored in R-exp_bad.
3.4.2 Augmenter
@@
3.4.3 Examiner (AI-synthetic data)
The ChatGPT-generated definitions underwent a second cleaning round using the same methodology as in Section 3.4.1 Examiner (expert-annotated data). Here, approximately 56% of the ChatGPT definitions were found suitable for model training, with the remaining being either contextually inappropriate or incompatible with the expert-provided lay definitions. The final tally was 96,668 ‘good’ and 75,175 ‘bad’ general definitions, stored in R-syn_good and R-syn_bad, respectively.
After the end of the EAE, we sampled 500 medical jargon terms each from the R-exp_good and R-syn_good datasets. We conducted human verification  [^7] The details about Data Quality Checking and Train/Eval/Test Split can be found in Appendix D on corresponding data points for these 1000 jargon terms. We obtained a high human agreement for both the R-exp_good and R-syn_good datasets’ quality. After correcting individual invalid data, we used this portion of the data as evaluation and test data (in a 1:1 ratio). We used the remaining R-exp_good and R-syn_good datasets as training data, ensuring that these 1,000 medical jargon terms would not appear in the training data. Table 5 shows the overall dataset statistics of all the README versions across the pipeline.
@@ 4.1 Automatic Evaluation Metrics
We evaluate the efficacy of our model in producing lay definitions by contrasting them with the ground-truth reference lay definitions, utilizing the ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005) metrics. However, they are based on the exact word overlap and therefore provide insight into the informativeness of the generated lay definitions but do not necessarily reflect their factual accuracy Maynez et al. (2020). We also employ Scispacy to extract medical concepts from the model-generated and the reference lay definitions. We then compute the F1 Scores for these concept lists, referred to as UMLS-F1, to specifically measure the factuality of the generated content  [^9] The details about UMLS-F1 can be found in Appendix E.
@@ 4.2 Experimental Setting
We use the following base models: GPT-2, DistilGPT2, BioGPT, and Llama2 in our experiments  [^10] More details about Experimental Setting can be found in Appendix F.. We use the following symbols:
@@
We use one jargon, ’EDG’, as an example and show the input prompt of different settings J2L, J+C2L, J+G2L, J+C+G2L on Table 7.
Our experiments were divided into five distinct parts. In Set-1, we aimed to evaluate the performance gap on the J2L task between open-source base models and GPT3.5/4 models. We do one-shot prompting for Llama2  [^11] Because smaller open source language models DistilGPT2/GPT2/BioGPT do not have the ability to complete instruction following under zero-shot or few-shot settings, we only compare the results of Llama2 as a representative of open source language models with GPT3.5/4 to see the gap., GPT-3.5-turbo, and GPT-4 with prompts in Table 6. Set-2 explored the varying data quality across different versions within our EAE pipeline, where we fine-tuned the base models on J+G2L task. Set-3 focused on evaluating the effects of various data selection strategies on data augmentation outcomes. To do this, we ranked AI-synthetic data (R-syn_good) using different methods in Section 3.5, selecting either the top N entries with the highest scores (e.g., ‘SEMANTIC’ in Table 2) or the bottom N entries with the lowest scores (e.g., ‘SEMANTIC_r’). These selections were then evenly combined with expert-annotated data (R-exp_good) at a one-to-one ratio. Following this, we fine-tuned the base models using these diverse, mixed datasets to determine the impact of each selection method on model performance. In Set-4, we investigated the effects of incorporating different types of information (EHR context and UMLS-retrieved general definition) into the model inputs. In Set-5, we fine-tuned models of varying sizes (ranging from DistilGPT2-88M to Llama2-7B) and compared the outcomes with GPT-3.5-turbo using best settings learned from previous Set-2, Set-3, and Set-4.
@@ 4.3 Results
We start with our experiments with some base models’ performance on the J2L dataset with Set-1. In Appendix D, we showed a high level of human evaluators’ agreement for the definitions generated by GPT-3.5-turbo, which underlines its efficacy in crafting human-readable explanations for biomedical concepts. Consequently, GPT-3.5-turbo serves as a strong baseline for quality in README lay definition generation, indicative of a standard that meets human satisfaction. Despite this, our analysis of Llama2, illustrated in Figure 3, reveals a significant performance gap. This discrepancy underscores the critical necessity for enhancing the capabilities of Llama2 and other open-source models to achieve high-quality output in lay definition generation tasks, thereby improving patient education together with systems like NoteAid. Additionally, we observed that GPT-3.5-turbo and GPT4 exhibited comparable proficiency in this task.
@@ Appendix C General Definition Retrieval and Preprocessing
Appendix D Data Quality Checking and Train/Eval/Test Split after EAE Pipeline
@@
Appendix E Factuality metrics: UMLS-F1

### v3

@@ Abstract
The advancement in healthcare has shifted focus toward patient-centric approaches, particularly in self-care and patient education, facilitated by access to Electronic Health Records (EHR). However, medical jargon in EHRs poses significant challenges in patient comprehension. To address this, we introduce a new task of automatically generating lay definitions, aiming to simplify complex medical terms into patient-friendly lay language. We first created the README dataset, an extensive collection of over 50,000 unique (medical term, lay definition) pairs and 300,000 mentions, each offering context-aware lay definitions manually annotated by domain experts. We have also engineered a data-centric Human-AI pipeline that synergizes data filtering, augmentation, and selection to improve data quality. We then used README as the training data for models and leveraged a Retrieval-Augmented Generation method to reduce hallucinations and improve the quality of model outputs. Our extensive automatic and human evaluations demonstrate that open-source mobile-friendly models, when fine-tuned with high-quality data, are capable of matching or even surpassing the performance of state-of-the-art closed-source large language models like ChatGPT. This research represents a significant stride in closing the knowledge gap in patient education and advancing patient-centric healthcare solutions.   [^1] Our code and data will be released at https://github.com/seasonyao/NoteAid-README
@@ 1 Introduction
Throughout the extensive history of natural language processing (NLP), enhancing individuals’ ability to read and comprehend complex texts has been a significant research goal that remains only partially achieved Zeng et al. (2020). Among the various tasks in this field, improving the comprehension of medical electronic health records (EHRs) stands out due to the complexity and specificity of medical terminology Nutbeam (2023). Efforts to enhance the comprehension of EHRs can not only advance NLP’s goal of aiding the understanding of complex texts but also hold substantial social significance by increasing efficiency and reducing errors for both patients and healthcare professionals Nanna et al. (2009); Boling (2009); Spooner et al. (2019); Baldry et al. (1986); Schillinger et al. (2009).
@@
However, despite advancements in EHRs management Walker et al. (2019), one significant barrier persists in the form of medical jargon   [^2] A “medical jargon term” is the specialized language healthcare professionals use that can be complex for non-medical individuals. A “lay definition” translates this jargon into accessible language for the general public, aiming to bridge the understanding gap and enhance the communication of health-related information. in EHRs, impeding patient understanding and self-care Kujala et al. (2022); Choudhry et al. (2016); Khasawneh et al. (2022). As shown in Figure 1, tools like NoteAid Chen et al. (2018); Kwon et al. (2022), which employs NLP to demystify complex medical terms, have been instrumental in bridging the communication gap between healthcare professionals and patients Lalor et al. (2018, 2021, 2023). However, lay language dictionary resources like the Consumer Health Vocabulary (CHV) Zeng and Tse (2006); He et al. (2017); Ibrahim et al. (2020) are limited in scale, posing a challenge to NoteAid. For instance, only a fraction (about 4%) of the medical terms in NoteAid have been annotated with lay definitions, highlighting the need for a more scalable solution to address this knowledge gap effectively.
@@
Addressing this issue requires shifting our focus to online health education materials such as Unified Medical Language System (UMLS)  Bodenreider (2004), MedlinePlus Patrias and Wendling (2007), Wikipedia, and Google. However, as indicated in Table 4, these resources often present too difficult information for the average patient to understand. For example, these resources’ average readability measured by the Flesch Kincaid Grade Level Solnyshkina et al. (2017) was post-secondary or higher education, while the average readability of a US adult was 7-8th grade level Doak et al. (1996, 1998); Eltorai et al. (2014). To bridge this gap, we have engaged medical experts to meticulously curate lay definitions for jargon terms found in NoteAid-MedJEx Kwon et al. (2022), targeting a comprehension level suitable for individuals with a 7th to 8th-grade education. Each jargon term has been redefined across various contexts, ensuring their applicability in diverse clinical scenarios. This effort led to our creation of the REsource of lAy Definitions for MEdical jargon (README) dataset, an expansive resource containing over 51,623 (medical jargon term, lay definition) pairs. The README dataset comprises an impressive 308,242 data points, each consisting of a clinical note context, a medical jargon term, and its corresponding lay definition. Thus, the dataset significantly enhances the accessibility and comprehensibility of medical information for patients.
@@
  • Introduced a new task of automatically generating lay definitions for medical jargon. We created a substantial expert-annotated dataset of 308,242 data points that can be used directly as a detailed lay-language dictionary for patient education tools such as NoteAid and as training data for this new task.
@@ 3.2 Lay Definition Annotation
Domain-experts read each sentence and identified as medical jargon terms that would be considered difficult to comprehend for anyone no greater than a 7th-grade education  [^3] The rule of thumb is that if a term has a lay definition comprehensible to a 4-7th grader as judged by FKGL Solnyshkina et al. (2017), this term is included as a jargon term.. Overall, 51,623 unique (medical jargon term, lay definition) pairs with 308,242 mentions  [^4] So there are 308,242 data points with (EHR context, medical jargon term, lay definition) format. in the EHR repository have been annotated by complying with the annotation guidelines presented in Appendix A.
@@ 3.3 General Definition Retrieval
We then employed the Scispacy library  [^5] We used the Scispacy en_core_sci_lg model to obtain the data. https://github.com/allenai/scispacy to retrieve the corresponding UMLS definitions Lindberg et al. (1993) of these annotated medical jargon terms as the general definitions. We follow the retrieval and preprocessing steps in Appendix C to filter the valid general definitions for the README dataset  [^6] We discuss the concept ambiguity issue in Appendix D. This preliminary cleaning results in 308,242 data points in the README-exp (e.g., expert annotated) dataset, each consisting of a clinical note context, a medical jargon term, a corresponding lay definition, and a corresponding general definition, as shown in Figure 2.
@@ 3.4 Examiner-Augmenter-Examiner (EAE)
Examiner (expert-annotated data)
Initially, basic data cleaning, as outlined in Appendix C, was applied. To enhance this, we employed GPT-3.5-turbo, using a few-shot learning approach with seven examples - four demonstrating acceptable data points and three showing unacceptable ones. These prompts served as the ‘Human’ element in our Human-AI-in-the-loop model, as depicted in Figure 2 and detailed in Algorithm 1. The prompts are detailed in Table 9. We choose GPT-3.5-turbo here because our evaluation (Section 4.3 and Appendix F) shows that the definitions it generates for medical terms can reach a human-satisfying level. Post-cleaning  [^7] More GPT Running Details are in Appendix B.4, approximately 39% of UMLS general definitions were deemed suitable by the Examiner (e.g., GPT-3.5-turbo). The suitable 113,659 data points were archived in R-exp_good, while the unsuitable 177,140 ones were stored in R-exp_bad.
Augmenter
@@
Examiner (AI-synthetic data)
The ChatGPT-generated definitions underwent a second cleaning round using the same methodology as in Section 3.4 Examiner (expert-annotated data). Here, approximately 56% of the ChatGPT definitions were found suitable for model training, with the remaining being either contextually inappropriate or incompatible with the expert-provided lay definitions. The final tally was 96,668 ‘good’ and 75,175 ‘bad’ general definitions, stored in R-syn_good and R-syn_bad, respectively. We discuss the EAE pipeline efficacy in Appendix E
Qaulty Checking
After the end of the EAE, we sampled 500 medical jargon terms each from the R-exp_good and R-syn_good datasets. We conducted human verification  [^8] The details about Data Quality Checking and Train/Eval/Test Split can be found in Appendix F on corresponding data points for these 1000 jargon terms. We obtained a high human agreement for both the R-exp_good and R-syn_good datasets’ quality. After correcting individual invalid data, we used this portion of the data as evaluation and test data (in a 1:1 ratio). We used the remaining R-exp_good and R-syn_good datasets as training data, ensuring that these 1,000 medical jargon terms would not appear in the training data. Table 5 shows the overall dataset statistics of all the README versions across the pipeline.
@@ 4.1 Automatic Evaluation Metrics
We evaluate the efficacy of our model in producing lay definitions by contrasting them with the ground-truth reference lay definitions, utilizing the ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005) metrics. However, they are based on the exact word overlap and therefore provide insight into the informativeness of the generated lay definitions but do not necessarily reflect their factual accuracy Maynez et al. (2020). We also employ Scispacy to extract medical concepts from the model-generated and the reference lay definitions. We then compute the F1 Scores for these concept lists, referred to as UMLS-F1, to specifically measure the factuality of the generated content  [^10] The details about UMLS-F1 can be found in Appendix G.
@@ 4.2 Experimental Setting
We use the following base models: GPT-2, DistilGPT2, BioGPT, and Llama2 in our experiments  [^11] More Experimental Setting details are in Appendix H.. We use the following symbols:
@@
We use one jargon, ’EDG’, as an example and show the input prompt of different settings J2L, J+C2L, J+G2L, J+C+G2L  [^12] We use ”context” and ”C” interchangeably to refer to the EHR context in which the jargon occurs. Similarly, we use ”general definition” and ”G” interchangeably to refer to the general definition retrieved from UMLS. on Table 8.
Our experiments were divided into five distinct parts. In Set-1, we aimed to evaluate the performance gap on the J2L task between open-source base models and GPT3.5/4 models. We do one-shot prompting for Llama2  [^13] Because smaller open source language models DistilGPT2/GPT2/BioGPT do not have the ability to complete instruction following under zero-shot or few-shot settings, we only compare the results of Llama2 as a representative of open source language models with GPT3.5/4 to see the gap., GPT-3.5-turbo, and GPT-4 with prompts in Table 6. Set-2 explored the varying data quality across different versions within our EAE pipeline, where we fine-tuned the base models on J+G2L task. Set-3 focused on evaluating the effects of various data selection strategies on data augmentation outcomes. To do this, we ranked AI-synthetic data (R-syn_good) using different methods in Section 3.5, selecting either the top N entries with the highest scores (e.g., ‘SEMANTIC’ in Table 2) or the bottom N entries with the lowest scores (e.g., ‘SEMANTIC_r’). These selections were then evenly combined with expert-annotated data (R-exp_good) at a one-to-one ratio. Following this, we fine-tuned the base models using these diverse, mixed datasets to determine the impact of each selection method on model performance. In Set-4, we investigated the effects of incorporating different types of information (EHR context and UMLS-retrieved general definition) into the model inputs. In Set-5, we fine-tuned models of varying sizes (ranging from DistilGPT2-88M to Llama2-7B) and compared the outcomes with GPT-3.5-turbo using best settings learned from previous Set-2, Set-3, and Set-4. More experiment designs and results can be found in Appendix I
@@ 4.3 Results
We start with our experiments with some base models’ performance on the J2L dataset with Set-1. In Appendix F, we showed a high level of human evaluators’ agreement for the definitions generated by GPT-3.5-turbo, which underlines its efficacy in crafting human-readable explanations for biomedical concepts. Consequently, GPT-3.5-turbo serves as a strong baseline for quality in README lay definition generation, indicative of a standard that meets human satisfaction. Despite this, our analysis of Llama2, illustrated in Figure 3, reveals a significant performance gap. This discrepancy underscores the critical necessity for enhancing the capabilities of Llama2 and other open-source models to achieve high-quality output in lay definition generation tasks, thereby improving patient education together with systems like NoteAid. Additionally, we observed that GPT-3.5-turbo and GPT4  [^14] All GPT experiments in our paper were conducted using Microsoft Azure, which can be used HIPAA-compliantly, ensuring the ethical handling of sensitive data. exhibited comparable proficiency in this task.
@@ Appendix C General Definition Retrieval and Preprocessing
Appendix D Discussion on Concept Ambiguity
An important aspect that merits discussion is the potential ambiguity of medical concepts and its impact on generating lay user explanations. The inherent complexity and context-dependency of medical terms can create challenges in crafting universally understandable definitions, potentially leading to patient misinterpretations. Our approach to addressing the ambiguity of medical concepts is rooted in the Word-Sense Disambiguation (WSD) phase, as elaborated in related works such as Kwon et al. (2022). The WSD phase links ambiguous medical terms to accurate, disambiguated definitions from medical concept dictionaries. These definitions serve as the foundation for generating lay definitions suitable for patient understanding, especially when dictionary definitions lack readability. In our study, the WSD phase is implicitly managed during the general definition retrieval stage using Scispacy and UMLS tools.
Appendix E Discussion on EAE Pipeline Efficacy
The primary objective of the EAE pipeline is to validate the effectiveness of Scispacy + UMLS for Word Sense Disambiguation (WSD), rather than addressing quality issues in the expert-annotated lay definitions. While quality checks and filtering for (jargon, context, lay definition) do not necessitate an LLM, they are essential for (jargon, context, lay definition, general definition) due to the noise introduced by Scispacy + UMLS tools. It is crucial to note that this noise is outside the scope of this paper.
An important aspect that merits discussion is the efficacy of the EAE pipeline design, particularly regarding the high number of data points categorized as R-exp_bad or R-syn_bad in Table 5. Ideally, a more efficient pipeline would achieve a higher number of R-exp_good data points, thereby reducing the need for additional verification rounds. However, our current settings prioritize reducing false positives, even if this results in an increased number of false negatives. Strict AI verification is essential to mitigate false positives, which could lead to patient misunderstandings if incorrect definitions are generated. Given the large size of the original README dataset, the harm of false negatives is more acceptable.
Therefore, we have intentionally set a stricter standard for AI during prompting (as shown in Table 9). This approach may classify some good-quality data as bad, but it ensures that any data passing the Examiner stage is of high quality. This is corroborated by the high human agreement rates observed in Appendix F. Consequently, the final small human verification step is manageable without significantly increasing the workload.
The EAE pipeline and related prompts effectively detect and filter (jargon, context, lay definition, general definition) with minimal human effort, ensuring a sufficient quantity of valid data for subsequent training. In scenarios where the dataset is smaller or where more false positives are acceptable, the sensitivity of the AI examiner may need to be adjusted. Nevertheless, the three stages of the E→A→E pipeline are crucial for maintaining data quality in our task, as highlighted in Table 1, and can be extended to other similar scenarios.
Appendix F Data Quality Checking and Train/Eval/Test Split after EAE Pipeline
@@
Appendix G Factuality metrics: UMLS-F1